OpenStudy (anonymous):

Solve for b,then find the value of b when a=3. 4a=2b-7 A. -9/2 B. 5/2 C. 19/2 D. 17 Please help

OpenStudy (anonymous):

\[4a=2b-7\] add \(7\)

OpenStudy (anonymous):

\[4a+7=2b\] divide by 2

OpenStudy (anonymous):

\[\frac{4a+7}{2}=b\]

OpenStudy (anonymous):

replace \(a\) by \(3\)